OpenStudy (anonymous):

Solve -2 I h-7 I=-28

OpenStudy (anonymous):

|h-7|=14 Do you know how to evaluate absolute value equations?

OpenStudy (anonymous):

(h-7) - 14 >= 0 (h-7) - 14 < 0

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

-(h-7) - 14 < 0 i mean

OpenStudy (anonymous):

h-7=14 or h-7=-14, solve for h in both equations

OpenStudy (anonymous):

here are the answer choices a. no solution b. h=21 c. h=-7, h=21 d. h=7, h=-21

OpenStudy (anonymous):

can you solve h-7=14?

OpenStudy (anonymous):

I got it I was looking at it wrong thanks